vb

GNU Info

for version 4.13, 18 September 2008

Brian J. Fox

This manual is for GNU Info (version 4.13, 18 September 2008), a program for viewing documents in Info format (usually created from Texinfo source files). Copyright \begin{matrix}c \\ ◯ \end{matrix} 1992, 1993, 1996, 1997, 1998, 1999, 2001, 2002, 2003, 2004, 2005, 2006, 2007, 2008 Free Software Foundation, Inc.

Permission is granted to copy, distribute and/or modify this document under the terms of the GNU Free Documentation License, Version 1.2 or any later version published by the Free Software Foundation; with no Invariant Sections, with the Front-Cover texts being “A GNU Manual”, and with the Back-Cover Texts as in (a) below. A copy of the license is included in the section entitled “GNU Free Documentation License’ in the Texinfo manual.

(a) The FSF’s Back-Cover Text is: “You have the freedom to copy and modify this GNU manual. Buying copies from the FSF supports it in developing GNU and promoting software freedom.”

This document is part of a collection distributed under the GNU Free Documentation License. If you want to distribute this document separately from the collection, you can do so by adding a copy of the license to the document, as described in section 6 of the license.

Table of Contents

1 Stand-alone Info . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 Invoking Info . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 3 Moving the Cursor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 4 Moving Text Within a Window . . . . . . . . . . . . . . . 8 5 Selecting a Node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 6 Searching an Info File . . . . . . . . . . . . . . . . . . . . . . . . . 14 7 Selecting Cross References . . . . . . . . . . . . . . . . . . . 16

7.1 Parts of an Xref. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 7.2 Selecting Xrefs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

8 Manipulating Multiple Windows . . . . . . . . . . . . . 18

8.1 The Mode Line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 8.2 Window Commands. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 8.3 The Echo Area. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

9 Printing Nodes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 10 Miscellaneous Commands . . . . . . . . . . . . . . . . . . . 24 11 Manipulating Variables . . . . . . . . . . . . . . . . . . . . . . 26 12 Customizing Key Bindings and Variables . . 29

12.1 Invoking infokey . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 12.2 infokey source format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

Appendix A Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32

1 Stand-alone Info

The Info program is a stand-alone program, part of the Texinfo distribution, which is used to view Info files on an ASCII terminal. Info files are typically the result of processing Texinfo files with the program makeinfo (also in the Texinfo distribution)

Texinfo itself is a documentation system that uses a single source file to produce both on-line information and printed output. You can typeset and print the files that you read in Info.

GNU Emacs also provides an Info reader (just type M-x info in Emacs). Emacs Info and stand-alone Info have nearly identical user interfaces, although customization and other details are different (this manual explains the stand-alone Info reader). The Emacs Info reader supports the X Window System and other such bitmapped interfaces, not just plain ASCII, so if you want the prettiest display for Info files, you should try it. You can use Emacs Info without using Emacs for anything else. (Type C-x C-c to exit; this also works in the stand-alone Info reader.)

2 Invoking Info

GNU Info accepts several options to control the initial node being viewed, and to specify which directories to search for Info files. Here is a template showing an invocation of GNU Info from the shell:

info [option]... [menu-item...]

The program accepts the following options:

--apropos=string
-k string

Specify a string to search in every index of every Info file installed on your system. Info looks up the named string in all the indices it can find, prints the results to standard output, and then exits. If you are not sure which Info file explains certain issues, this option is your friend. (If your system has a lot of Info files installed, searching all of them might take some time!) You can invoke the apropos command from inside Info; see Chapter 6 [Searching Commands], page 14.

--directory directory-path
-d directory-path

Prepend directory-path to the list of directory paths searched when Info needs to find a file. You may issue --directory multiple times; once for each directory which contains Info files. The list of directories searched by Info is constructed from the value of the environment variable INFOPATH; --directory causes the named directory-path to be prepended to that list. The value of INFOPATH is a list of directories usually separated by a colon; on MS-DOS/MS-Windows systems, the semicolon is used. If you do not define INFOPATH, Info uses a default path defined when Info was built as the initial list of directories. If the value of INFOPATH ends with a colon (or semicolon on MS-DOS/MS-Windows), the initial list of directories is constructed by appending the build-time default to the value of INFOPATH.

--dribble=dribble-file

Specify a file where all user keystrokes will be recorded. This file can be used later to replay the same sequence of commands, see the ‘--restore’ option below.

--file filename
-f filename

Specify a particular Info file to visit. By default, Info visits the file dir; if you use this option, Info will start with (filename)Top as the first file and node. If filename is an absolute file name, or begins with ‘./’ or ‘../’, Info looks for filename only in the directory of the specified filename, and adds the directory of filename to the value of INFOPATH. In contrast, if filename is in the form of a relative file name, but without the ‘./’ or ‘../’ prefix, Info will only look for it in the directories specified in INFOPATH. In other words, Info does not treat file names which lack ‘./’ and ‘../’ prefix as relative to the current directory.

In every directory Info tries, if filename is not found, Info looks for it with a number of known extensions of Info files^{1}. For every known extension, Info looks for a compressed file, if a regular file isn’t found. Info supports files compressed

with gzip, bzip2, compress and yabba programs; it calls gunzip, bunzip2,

uncompress and unyabba, accordingly, to decompress such files. Compressed Info files are assumed to have ‘.z’, ‘.gz’, ‘.bz2’, ‘.Z’, or ‘.Y’ extensions, possibly in addition to one of the known Info files extensions^{2}.

--help

-h Produces a relatively brief description of the available Info options.

--index-search string

After processing all command-line arguments, go to the index in the Info file and search for index entries which match string. If such an entry is found, the Info session begins with displaying the node pointed to by the first matching index entry; press , to step through the rest of the matching entries. If no such entry exists, print ‘no entries found’ and exit with nonzero status. This can be used from another program as a way to provide online help, or as a quick way of starting to read an Info file at a certain node when you don’t know the exact name of that node. This command can also be invoked from inside Info; see Chapter 6 [Searching Commands], page 14.

--node nodename
-n nodename

Specify a particular node to visit in the initial file that Info loads. This is especially useful in conjunction with --file^{3}. You may specify --node multiple times; for an interactive Info, each nodename is visited in its own window, for a non-interactive Info (such as when --output is given) each nodename is processed sequentially.

--output filename
-o filename

Specify filename as the name of a file to which to direct output. Each node that Info visits will be output to filename instead of interactively viewed. A value of - for filename specifies the standard output.

--raw-escapes
--no-raw-escapes

-R Do not remove ANSI escape sequences from documents. Some versions of Groff, the GNU document formatter, produce man pages with ANSI escape sequences for bold, italics, and underlined characters, and for colorized text. By default, Info lets those escape sequences pass through directly to the terminal. If your terminal does not support these escapes, use --no-raw-escapes to make Info remove them.

1 ‘.info’, ‘-info’, ‘/index’, and ‘.inf’.

2 The MS-DOS version allows for the Info extension, such as .inf, and the short compressed file extensions,

such as ‘.z’ and ‘.gz’, to be merged into a single extension, since DOS doesn’t allow more than a single dot in the basename of a file. Thus, on MS-DOS, if Info looks for ‘bison’, file names like ‘bison.igz’ and ‘bison.inz’ will be found and decompressed by gunzip.

3 Of course, you can specify both the file and node in a --node command; but don’t forget to escape the

open and close parentheses and whitespace from the shell as in: info --node "(emacs)Buffers".

--restore=dribble-file

Read keystrokes from dribble-file, presumably recorded during previous Info session (see the description of the ‘--dribble’ option above). When the keystrokes in the files are all read, Info reverts its input to the usual interactive operation.

--show-malformed-multibytes
--no-show-malformed-multibytes

Show malformed multibyte sequences in the output. By default, such sequences are dropped.

--show-options
--usage

-O This option causes Info to look for the node that describes how to invoke the program and its command-line options, and begin the session by displaying that node. It is provided to make it easier to find the most important usage information in a manual without the need to wade through complex menu hierarchies. The effect is similar to the M-x goto-invocation command (see [goto-invocation], page 12) from inside Info.

--speech-friendly

-b On MS-DOS/MS-Windows only, this option causes Info to use standard file I/O functions for screen writes. (By default, Info uses direct writes to the video memory on these systems, for faster operation and colored display support.) This allows the speech synthesizers used by blind persons to catch the output and convert it to audible speech.

--subnodes

This option only has meaning when given in conjunction with --output. It means to recursively output the nodes appearing in the menus of each node being output. Menu items which resolve to external Info files are not output, and neither are menu items which are members of an index. Each node is only output once.

--version

Prints the version information of Info and exits.

--vi-keys

This option binds functions to keys differently, to emulate the key bindings of vi and Less. The default key bindings are generally modeled after Emacs. (See Chapter 12 [Custom Key Bindings], page 29, for a more general way of altering GNU Info’s key bindings.)

--where
--location

-w Show the filename that would be read and exit, instead of actually reading it and starting Info.

menu-item

Info treats its remaining arguments as the names of menu items. The first argument is a menu item in the initial node visited (generally dir), the second argument is a menu item in the first argument’s node, etc. You can easily move to the node of your choice by specifying the menu names which describe the path to that node. For example,

               info emacs buffers

first selects the menu item ‘Emacs’ in the node ‘(dir)Top’, and then selects the menu item ‘Buffers’ in the node ‘(emacs)Top’.

To avoid searching the ‘dir’ files and just show some arbitrary file, use ‘-f’ and the

filename, as in ‘info -f ./foo.info’.

The index search and the search for the node which describes program invocation and command-line options begins after processing all the command-line menu items. Therefore, the Info file searched for the index or the invocation node is the file where Info finds itself after following all the menu items given on the command line. This is so ‘info emacs --show-options’ does what you’d expect.

Finally, Info defines many default key bindings and variables. See Chapter 12 [Custom Key Bindings], page 29, for information on how to customize these settings.

3 Moving the Cursor

Many people find that reading screens of text page by page is made easier when one is able to indicate particular pieces of text with some kind of pointing device. Since this is the case, GNU Info (both the Emacs and stand-alone versions) have several commands which allow you to move the cursor about the screen. The notation used in this manual to describe keystrokes is identical to the notation used within the Emacs manual, and the GNU Readline manual. See Section “User Input” in The GNU Emacs Manual, if you are unfamiliar with the notation.^{1}

The following table lists the basic cursor movement commands in Info. Each entry consists of the key sequence you should type to execute the cursor movement, the M-x^{2} command name (displayed in parentheses), and a short description of what the command does. All of the cursor motion commands can take a numeric argument (see Chapter 10 [to find out how to supply them], page 24. With a numeric argument, the motion commands are simply executed that many times; for example, a numeric argument of 4 given to next-line causes the cursor to move down 4 lines. With a negative numeric argument, the motion is reversed; an argument of -4 given to the next-line command would cause the cursor to move up 4 lines.

C-n (next-line)

DOWN (an arrow key)

Move the cursor down to the next line.

C-p (prev-line)

UP (an arrow key)

Move the cursor up to the previous line.

C-a (beginning-of-line)

HOME (on DOS/Windows only)

Move the cursor to the start of the current line.

C-e (end-of-line)

END (on DOS/Windows only)

Move the cursor to the end of the current line.

C-f (forward-char)

RIGHT (an arrow key)

Move the cursor forward a character.

C-b (backward-char)

LEFT (an arrow key)

Move the cursor backward a character.

M-f (forward-word)

C-RIGHT (on DOS/Windows only)

Move the cursor forward a word.

1 Here’s a short summary. C-x means press the CTRL key and the key x. M-x means press the META key

and the key x. On many terminals th META key is known as the ALT key. SPC is the space bar. The other

keys are usually called by the names imprinted on them.

2 M-x is also a command; it invokes execute-extended-command. See Section “Executing an extended

command” in The GNU Emacs Manual, for more detailed information.

M-b (backward-word)

C-LEFT (on DOS/Windows only)

Move the cursor backward a word.

M-< (beginning-of-node)

C-HOME (on DOS/Windows only)

b

M-b, vi-like operation Move the cursor to the start of the current node.

M-> (end-of-node)

C-END (on DOS/Windows only) e Move the cursor to the end of the current node.

M-r (move-to-window-line)

Move the cursor to a specific line of the window. Without a numeric argument, M-r moves the cursor to the start of the line in the center of the window. With a numeric argument of n, M-r moves the cursor to the start of the nth line in the window.

4 Moving Text Within a Window

Sometimes you are looking at a screenful of text, and only part of the current paragraph you are reading is visible on the screen. The commands detailed in this section are used to shift which part of the current node is visible on the screen.

Scrolling commands are bound differently when ‘--vi-keys’ operation is in effect (see [–vi-keys], page 4). These key bindings are designated with “vi-like operation”. See Chapter 12 [Custom Key Bindings], page 29, for information on arbitrarily customizing key bindings and variable settings.

SPC (scroll-forward)

Shift the text in this window up. That is, show more of the node which is currently below the bottom of the window. With a numeric argument, show that many more lines at the bottom of the window; a numeric argument of 4 would shift all of the text in the window up 4 lines (discarding the top 4 lines), and show you four new lines at the bottom of the window. Without a numeric argument, SPC takes the bottom two lines of the window and places them at the top of the window, redisplaying almost a completely new screenful of lines. If you are at the end of a node, SPC takes you to the “next” node, so that you can read an entire manual from start to finish by repeating SPC.

The default scroll size is one screen-full, but it can be changed by invoking the

(scroll-forward-page-only-set-window) command, ‘z’ under ‘--vi-keys’,

with a numeric argument.

NEXT (an arrow key) (scroll-forward-page-only)
C-v

C-f, vi-like operation f, vi-like operation M-SPC, vi-like operation

Shift the text in this window up. This is identical to the SPC operation above, except that it never scrolls beyond the end of the current node.

The NEXT key is known as the PAGEDOWN key on some keyboards.

z (scroll-forward-page-only-set-window, vi-like operation)

Scroll forward, like with NEXT, but if a numeric argument is specified, it becomes the default scroll size for subsequent scroll-forward and scrollbackward commands and their ilk.

DEL (scroll-backward)

Shift the text in this window down. The inverse of scroll-forward. If you are at the start of a node, DEL takes you to the “previous” node, so that you can read an entire manual from finish to start by repeating DEL. The default scroll size can be changed by invoking the (scroll-backward-page-only-set-window) command, ‘w’ under ‘--vi-keys’, with a numeric argument.

PREVIOUS (arrow key) (scroll-backward-page-only)

PRIOR (arrow key)

M-v

b, vi-like operation C-b, vi-like operation

Shift the text in this window down. The inverse of scroll-forward-page-only. Does not scroll beyond the start of the current node. The default scroll size can be changed by invoking the(scroll-backward-page-only-set-window) command, ‘w’ under ‘--vi-keys’, with a numeric argument.

w (scroll-backward-page-only-set-window, vi-like operation)

Scroll backward, like with PREVIOUS, but if a numeric argument is specified, it becomes the default scroll size for subsequent scroll-forward and scrollbackward commands.

C-n (down-line, vi-like operation) C-e, vi-like operation RET, vi-like operation LFD, vi-like operation DOWN, vi-like operation

Scroll forward by one line. With a numeric argument, scroll forward that many lines.

C-p (up-line, vi-like operation) UP, vi-like operation y, vi-like operation k, vi-like operation C-k, vi-like operation C-y, vi-like operation

Scroll backward one line. With a numeric argument, scroll backward that many lines.

d (scroll-half-screen-down, vi-like operation)

C-d, vi-like operation Scroll forward by half of the screen size. With a numeric argument, scroll that many lines. If an argument is specified, it becomes the new default number of lines to scroll for subsequent ‘d’ and ‘u’ commands.

u (scroll-half-screen-up, vi-like operation)

C-u, vi-like operation Scroll back by half of the screen size. With a numeric argument, scroll that many lines. If an argument is specified, it becomes the new default number of lines to scroll for subsequent ‘u’ and ‘d’ commands.

The scroll-forward and scroll-backward commands can also move forward and backward through the node structure of the file. If you press SPC while viewing the end of a node, or DEL while viewing the beginning of a node, what happens is controlled by the variable scroll-behavior (see [scroll-behavior], page 27).

The scroll-forward-page-only and scroll-backward-page-only commands never

scroll beyond the current node.

The PREVIOUS key is the PAGEUP key on many keyboards. Emacs refers to it by the name PRIOR. When you use PRIOR or PAGEUP to scroll, Info never scrolls beyond the beginning of the current node.

If your keyboard lacks the DEL key, look for a key called BS, or ‘Backspace’, sometimes designated with an arrow which points to the left, which should perform the same function.

C-l (redraw-display)

Redraw the display from scratch, or shift the line containing the cursor to a specified location. With no numeric argument, ‘C-l’ clears the screen, and then redraws its entire contents. Given a numeric argument of n, the line containing the cursor is shifted so that it is on the nth line of the window.

C-x w (toggle-wrap)

Toggles the state of line wrapping in the current window. Normally, lines which are longer than the screen width wrap, i.e., they are continued on the next line. Lines which wrap have a ‘\’ appearing in the rightmost column of the screen. You can cause such lines to be terminated at the rightmost column by changing the state of line wrapping in the window with C-x w. When a line which needs more space than one screen width to display is displayed, a ‘$’ appears in the rightmost column of the screen, and the remainder of the line is invisible. When long lines are truncated, the modeline displays the ‘$’ character near its left edge.

5 Selecting a Node

This section details the numerous Info commands which select a new node to view in the current window.

The most basic node commands are ‘n’, ‘p’, ‘u’, and ‘l’. Note that the commands to select nodes are mapped differently when ‘--vi-keys’ is in effect; these keybindings are designated below as “vi-like operation”.

When you are viewing a node, the top line of the node contains some Info pointers which describe where the next, previous, and up nodes are. Info uses this line to move about the node structure of the file when you use the following commands:

n (next-node)

C-NEXT (on DOS/Windows only) C-x n, vi-like operation

Select the ‘Next’ node.

The NEXT key is known as the PGDN key on some keyboards.

p (prev-node)

C-PREVIOUS (on DOS/Windows only)

Select the ‘Prev’ node.

The PREVIOUS key is known as the PGUP key on some keyboards.

u (up-node)

C-UP (an arrow key on DOS/Windows only) C-x u, vi-like operation

Select the ‘Up’ node.

You can easily select a node that you have already viewed in this window by using the ‘l’ command—this name stands for “last”, and actually moves backwards through the history of visited nodes for this window. This is handy when you followed a reference to another node, possibly to read about a related issue, and would like then to resume reading at the same place where you started the excursion.

Each node where you press ‘l’ is discarded from the history. Thus, by the time you get to the first node you visited in a window, the entire history of that window is discarded.

l (history-node)

C-CENTER (on DOS/Windows only) ', vi-like operation

Pop the most recently selected node in this window from the node history.

Two additional commands make it easy to select the most commonly selected nodes; they are ‘t’ and ‘d’.

t (top-node)

M-t, vi-like operation Select the node ‘Top’ in the current Info file.

d (dir-node)

M-d, vi-like operation Select the directory node (i.e., the node ‘(dir)’).

Here are some other commands which immediately result in the selection of a different node in the current window:

< (first-node)

g, vi-like operation Selects the first node which appears in this file. This node is most often ‘Top’, but it does not have to be. With a numeric argument N, select the Nth node (the first node is node 1). An argument of zero is the same as the argument of 1.

> (last-node)

G, vi-like operation Select the last node which appears in this file. With a numeric argument N, select the Nth node (the first node is node 1). An argument of zero is the same as no argument, i.e., it selects the last node.

] (global-next-node)

Move forward or down through node structure. If the node that you are currently viewing has a ‘Next’ pointer, that node is selected. Otherwise, if this node has a menu, the first menu item is selected. If there is no ‘Next’ and no menu, the same process is tried with the ‘Up’ node of this node.

[ (global-prev-node)

Move backward or up through node structure. If the node that you are currently viewing has a ‘Prev’ pointer, that node is selected. Otherwise, if the node has an ‘Up’ pointer, that node is selected, and if it has a menu, the last item in the menu is selected.

You can get the same behavior as global-next-node and global-prev-node while simply scrolling through the file with SPC and DEL (see [scroll-behavior], page 27).

g (goto-node)

C-x g, vi-like operation Read the name of a node and select it. While reading the node name, completion (see Section 8.3 [The Echo Area], page 19) is only done for the nodes which reside in one of the Info files that were loaded in the current Info session; if the desired node resides in some other file, you must type the node exactly as it appears in that Info file, and you must include the Info file of the other file. For example,

     g(emacs)Buffers

finds the node ‘Buffers’ in the Info file ‘emacs’.

O (goto-invocation

I Read the name of a program and look for a node in the current Info file which describes the invocation and the command-line options for that program. The default program name is derived from the name of the current Info file. This command does the same as the ‘--show-options’ command-line option (see [–show-options], page 4), but it also allows to specify the program name; this is important for those manuals which describe several programs. If you need to find the Invocation node of a program that is documented in another Info file, you need to visit that file before invoking ‘I’. For example, if you are reading the Emacs manual and want to see the command-line options

of the makeinfo program, type g (texinfo) RET and then I makeinfo RET.

If you don’t know what Info file documents the command, or if invoking ‘I’ doesn’t display the right node, go to the ‘(dir)’ node (using the ‘d’ command) and invoke ‘I’ from there.

G (menu-sequence)

Read a sequence of menu entries and follow it. Info prompts for a sequence of menu items separated by commas. (Since commas are not allowed in a node name, they are a natural choice for a delimiter in a list of menu items.) Info then looks up the first item in the menu of the node ‘(dir)’ (if the ‘(dir)’ node cannot be found, Info uses ‘Top’). If such an entry is found, Info goes to the node it points to and looks up the second item in the menu of that node, etc. In other words, you can specify a complete path which descends through the menu hierarchy of a particular Info file starting at the ‘(dir)’ node. This has the same effect as if you typed the menu item sequence on Info’s command line, see [Info command-line arguments processing], page 4. For example,

      G Texinfo,Overview,Reporting Bugs RET

displays the node ‘Reporting Bugs’ in the Texinfo manual. (You don’t actually need to type the menu items in their full length, or in their exact letter-case. However, if you do type the menu items exactly, Info will find it faster.) If any of the menu items you type are not found, Info stops at the last entry it did find and reports an error.

C-x k (kill-node)

Kill a node. The node name is prompted for in the echo area, with a default of the current node. Killing a node means that Info tries hard to forget about it, removing it from the list of history nodes kept for the window where that node is found. Another node is selected in the window which contained the killed node.

C-x C-f (view-file)

Read the name of a file and selects the entire file. The command

     C-x C-f filename

is equivalent to typing

     g(filename)*
C-x C-b (list-visited-nodes)

Make a window containing a menu of all of the currently visited nodes. This window becomes the selected window, and you may use the standard Info commands within it.

C-x b (select-visited-node)

Select a node which has been previously visited in a visible window. This is similar to ‘C-x C-b’ followed by ‘m’, but no window is created.

6 Searching an Info File

GNU Info allows you to search for a sequence of characters throughout an entire Info file, search through the indices of an Info file, or find areas within an Info file which discuss a particular topic.

s (search)

/ Read a string in the echo area and search for it, either as a regular expression (by default) or a literal string. If the string includes upper-case characters, the Info file is searched case-sensitively; otherwise Info ignores the letter case. With a numeric argument of N, search for Nth occurrence of the string. Negative arguments search backwards.

? (search-backward, vi-like operation)

Read a string in the echo area and search backward through the Info file for that string. If the string includes upper-case characters, the Info file is searched casesensitively; otherwise Info ignores the letter case. With a numeric argument of N, search for Nth occurrence of the string. Negative arguments search forward.

R (toggle-regexp)

Toggle between using regular expressions and literal strings for searching. Info uses so-called ‘extended’ regular expression syntax, similar to Emacs (see Section “Using Regular Expressions” in The GNU Emacs Manual).

S (search-case-sensitively

Read a string in the echo area and search for it case-sensitively, even if the string includes only lower-case letters. With a numeric argument of N, search for Nth occurrence of the string. Negative arguments search backwards.

C-x n (search-next)

n, vi-like operation Search for the same string used in the last search command, in the same direction, and with the same case-sensitivity option. With a numeric argument of N, search for Nth next occurrence.

C-x N (search-previous)

N, vi-like operation Search for the same string used in the last search command, and with the same case-sensitivity option, but in the reverse direction. With a numeric argument of N, search for Nth previous occurrence.

C-s (isearch-forward)

Interactively search forward through the Info file for a string as you type it. If the string includes upper-case characters, the search is case-sensitive; otherwise Info ignores the letter case.

C-r (isearch-backward)

Interactively search backward through the Info file for a string as you type it. If the string includes upper-case characters, the search is case-sensitive; otherwise Info ignores the letter case.

i (index-search)

Look up a string in the indices for this Info file, and select a node to which the found index entry points.

, (next-index-match)

Move to the node containing the next matching index item from the last ‘i’ command.

M-x index-apropos

Grovel the indices of all the known Info files on your system for a string, and build a menu of the possible matches.

The most basic searching command is ‘s’ or ‘/’ (search). The ‘s’ command prompts you for a string in the echo area, and then searches the remainder of the Info file for an occurrence of that string. If the string is found, the node containing it is selected, and the cursor is left positioned at the start of the found string. Subsequent ‘s’ commands show you the default search string within ‘[’ and ‘]’; pressing RET instead of typing a new string will use the default search string. Under ‘--vi-keys’ (see [–vi-keys], page 4), using the ‘n’ or ‘N’ commands is a faster way of searching for the same string.

Incremental searching is similar to basic searching, but the string is looked up while you are typing it, instead of waiting until the entire search string has been specified.

Both incremental and non-incremental search by default ignore the case of letters when comparing the Info file text with the search string. However, an uppercase letter in the search string makes the search case-sensitive. You can force a case-sensitive non-incremental search, even for a string that includes only lower-case letters, by using the ‘S’ command (search-case-sensitively). The ‘n’ and ‘N’ commands operate case-sensitively if the last search command was ‘S’.

The most efficient means of finding something quickly in a manual is the ‘i’ command (index-search). This command prompts for a string, and then looks for that string in all the indices of the current Info manual. If it finds a matching index entry, it displays the node to which that entry refers and prints the full text of the entry in the echo area. You can press ‘,’ (next-index-match) to find more matches. A good Info manual has all of its important concepts indexed, so the ‘i’ command lets you use a manual as a reference.

If you don’t know what manual documents something, try the M-x index-apropos command. It prompts for a string and then looks up that string in all the indices of all the Info documents installed on your system. It can also be invoked from the command line; see [–apropos], page 2.

7 Selecting Cross References

We have already discussed the ‘Next’, ‘Prev’, and ‘Up’ pointers which appear at the top of a node. In addition to these pointers, a node may contain other pointers which refer you to a different node, perhaps in another Info file. Such pointers are called cross references, or xrefs for short.

7.1 Parts of an Xref

Cross references have two major parts: the first part is called the label; it is the name that you can use to refer to the cross reference, and the second is the target; it is the full name of the node that the cross reference points to.

The target is separated from the label by a colon ‘:’; first the label appears, and then the target. For example, in the sample menu cross reference below, the single colon separates the label from the target.

     * Foo Label: Foo Target.        More information about Foo.

Note the ‘.’ which ends the name of the target. The ‘.’ is not part of the target; it serves only to let Info know where the target name ends.

A shorthand way of specifying references allows two adjacent colons to stand for a target name which is the same as the label name:

     * Foo Commands::                Commands pertaining to Foo.

In the above example, the name of the target is the same as the name of the label, in

this case Foo Commands.

You will normally see two types of cross reference while viewing nodes: menu references, and note references. Menu references appear within a node’s menu; they begin with a ‘*’ at the beginning of a line, and continue with a label, a target, and a comment which describes what the contents of the node pointed to contains.

Note references appear within the body of the node text; they begin with *Note, and continue with a label and a target.

Like ‘Next’, ‘Prev’, and ‘Up’ pointers, cross references can point to any valid node. They are used to refer you to a place where more detailed information can be found on a particular subject. Here is a cross reference which points to a node within the Texinfo documentation: See Section “Writing an Xref” in the Texinfo Manual, for more information on creating your own texinfo cross references.

7.2 Selecting Xrefs

The following table lists the Info commands which operate on menu items.

1 (menu-digit)

2 … 9 M-1, vi-like operation M-2 … M-9, vi-like operation

Within an Info window, pressing a single digit, (such as ‘1’), selects that menu item, and places its node in the current window. For convenience, there is one exception; pressing ‘0’ selects the last item in the node’s menu. When ‘--vi-keys’ is in effect, digits set the numeric argument, so these commands are remapped to their ‘M-’ varieties. For example, to select the last menu item, press M-0.

0 (last-menu-item)

M-0, vi-like operation Select the last item in the current node’s menu.

m (menu-item)

Reads the name of a menu item in the echo area and selects its node. Completion is available while reading the menu label. See Section 8.3 [The Echo Area], page 19.

M-x find-menu

Move the cursor to the start of this node’s menu.

This table lists the Info commands which operate on cross references.

f (xref-item)
r

M-f, vi-like operation C-x r, vi-like operation

Reads the name of a note cross reference in the echo area and selects its node. Completion is available while reading the cross reference label. See Section 8.3 [The Echo Area], page 19.

Finally, the next few commands operate on menu or note references alike:

TAB (move-to-next-xref)

Move the cursor to the start of the next nearest menu item or note reference in this node. You can then use RET (select-reference-this-line) to select the menu or note reference.

M-TAB (move-to-prev-xref)

Shift-TAB (on DOS/Windows only)

Move the cursor the start of the nearest previous menu item or note reference in this node. On DOS/Windows only, the Shift-TAB key is an alias for M-TAB. This key is sometimes called ‘BackTab’.

RET (select-reference-this-line)

M-g, vi-like operation Select the menu item or note reference appearing on this line.

8 Manipulating Multiple Windows

A window is a place to show the text of a node. Windows have a view area where the text of the node is displayed, and an associated mode line, which briefly describes the node being viewed.

GNU Info supports multiple windows appearing in a single screen; each window is separated from the next by its modeline. At any time, there is only one active window, that is, the window in which the cursor appears. There are commands available for creating windows, changing the size of windows, selecting which window is active, and for deleting windows.

8.1 The Mode Line

A mode line is a line of inverse video which appears at the bottom of an Info window. It describes the contents of the window just above it; this information includes the name of the file and node appearing in that window, the number of screen lines it takes to display the node, and the percentage of text that is above the top of the window. It can also tell you if the indirect tags table for this Info file needs to be updated, and whether or not the Info file was compressed when stored on disk.

Here is a sample mode line for a window containing an uncompressed file named ‘dir’, showing the node ‘Top’.

     -----Info: (dir)Top, 40 lines --Top-------------------------------------
                 ^^^^^^^^
               (file)Node #lines    where

When a node comes from a file which is compressed on disk, this is indicated in the mode line with two small ‘z’’s. In addition, if the Info file containing the node has been split into subfiles, the name of the subfile containing the node appears in the modeline as well:

     --zz-Info: (emacs)Top, 291 lines --Top-- Subfile: emacs-1.Z-------------

Truncation of long lines (as opposed to wrapping them to the next display line, see Chapter 4 [Scrolling Commands], page 8) is indicated by a ‘$’ at the left edge of the mode line:

     --$--Info: (texinfo)Top, 480 lines --Top-- Subfile: texinfo-1-----------

When Info makes a node internally, such that there is no corresponding info file on disk, the name of the node is surrounded by asterisks (‘*’). The name itself tells you what the contents of the window are; the sample mode line below shows an internally constructed node showing possible completions:

     -----Info: *Completions*, 7 lines --All---------------------------------

8.2 Window Commands

It can be convenient to view more than one node at a time. To allow this, Info can display more than one window. Each window has its own mode line (see Section 8.1 [The Mode Line], page 18) and history of nodes viewed in that window (see Chapter 5 [history-node], page 11).

C-x o (next-window)

Select the next window on the screen. Note that the echo area can only be selected if it is already in use, and you have left it temporarily. Normally, ‘C-x o’ simply moves the cursor into the next window on the screen, or if you are already within the last window, into the first window on the screen. Given a numeric argument, ‘C-x o’ moves over that many windows. A negative argument causes ‘C-x o’ to select the previous window on the screen.

M-x prev-window

Select the previous window on the screen. This is identical to ‘C-x o’ with a negative argument.

C-x 2 (split-window)

Split the current window into two windows, both showing the same node. Each window is one half the size of the original window, and the cursor remains in the original window. The variable automatic-tiling can cause all of the windows on the screen to be resized for you automatically (see Chapter 11 [automatic-tiling], page 26).

C-x 0 (delete-window)

Delete the current window from the screen. If you have made too many windows and your screen appears cluttered, this is the way to get rid of some of them.

C-x 1 (keep-one-window)

Delete all of the windows excepting the current one.

ESC C-v (scroll-other-window)

Scroll the other window, in the same fashion that ‘C-v’ might scroll the current window. Given a negative argument, scroll the “other” window backward.

C-x ^(grow-window)

Grow (or shrink) the current window. Given a numeric argument, grow the current window that many lines; with a negative numeric argument, shrink the window instead.

C-x t (tile-windows)

Divide the available screen space among all of the visible windows. Each window is given an equal portion of the screen in which to display its contents. The variable automatic-tiling can cause tile-windows to be called when a window is created or deleted. See Chapter 11 [automatic-tiling], page 26.

8.3 The Echo Area

The echo area is a one line window which appears at the bottom of the screen. It is used to display informative or error messages, and to read lines of input from you when that is necessary. Almost all of the commands available in the echo area are identical to their Emacs counterparts, so please refer to that documentation for greater depth of discussion on the concepts of editing a line of text. The following table briefly lists the commands that are available while input is being read in the echo area:

C-f (echo-area-forward)

RIGHT (an arrow key) M-h, vi-like operation

Move forward a character.

C-b (echo-area-backward)

LEFT (an arrow key) M-l, vi-like operation

Move backward a character.

C-a (echo-area-beg-of-line)

M-0, vi-like operation Move to the start of the input line.

C-e (echo-area-end-of-line)

M-$, vi-like operation Move to the end of the input line.

M-f (echo-area-forward-word)

C-RIGHT (DOS/Windows only) M-w, vi-like operation

Move forward a word.

On DOS/Windows, C-RIGHT moves forward by words.

M-b (echo-area-backward-word)

C-LEFT (DOS/Windows only)

Move backward a word.

On DOS/Windows, C-LEFT moves backward by words.

C-d (echo-area-delete)

M-x, vi-like operation Delete the character under the cursor.

DEL (echo-area-rubout)

Delete the character behind the cursor.

On some keyboards, this key is designated BS, for ‘Backspace’. Those keyboards will usually bind DEL in the echo area to echo-area-delete.

C-g (echo-area-abort)

C-u, vi-like operation Cancel or quit the current operation. If completion is being read, this command discards the text of the input line which does not match any completion. If the input line is empty, it aborts the calling function.

RET (echo-area-newline)

Accept (or forces completion of) the current input line.

C-q (echo-area-quoted-insert)

C-v, vi-like operation Insert the next character verbatim. This is how you can insert control characters into a search string, for example, or the ‘?’ character when Info prompts with completion. printing character (echo-area-insert)

Insert the character. Characters that have their 8th bit set, and not bound to ‘M-’ commands, are also inserted verbatim; this is useful for terminals which support Latin scripts.

M-TAB (echo-area-tab-insert)

Shift-TAB (on DOS/Windows only)

Insert a TAB character.

On DOS/Windows only, the Shift-TAB key is an alias for M-TAB. This key is sometimes called ‘BackTab’.

C-t (echo-area-transpose-chars)

Transpose the characters at the cursor.

The next group of commands deal with killing, and yanking text^{1}. For an in-depth discussion of killing and yanking, see Section “Killing and Deleting” in the GNU Emacs Manual.

M-d (echo-area-kill-word)

M-X, vi-like operation Kill the word following the cursor.

M-DEL (echo-area-backward-kill-word)

M-BS Kill the word preceding the cursor.

On some keyboards, the ‘Backspace’ key is used instead of DEL, so M-BACKSPACE has the same effect as M-DEL.

C-k (echo-area-kill-line)

Kill the text from the cursor to the end of the line.

C-x DEL (echo-area-backward-kill-line)

Kill the text from the cursor to the beginning of the line.

C-y (echo-area-yank)

Yank back the contents of the last kill.

M-y (echo-area-yank-pop)

Yank back a previous kill, removing the last yanked text first.

Sometimes when reading input in the echo area, the command that needed input will only accept one of a list of several choices. The choices represent the possible completions, and you must respond with one of them. Since there are a limited number of responses you can make, Info allows you to abbreviate what you type, only typing as much of the response as is necessary to uniquely identify it. In addition, you can request Info to fill in as much of the response as is possible; this is called completion.

The following commands are available when completing in the echo area:

TAB (echo-area-complete)

SPC Insert as much of a completion as is possible.

1 Some people are used to calling these operations cut and paste, respectively.

? (echo-area-possible-completions)

Display a window containing a list of the possible completions of what you have typed so far. For example, if the available choices are:

     bar foliate food forget

and you have typed an ‘f’, followed by ‘?’, Info will pop up a window showing a node called ‘*Completions*’ which lists the possible completions like this:

     3 completions:
     foliate         food
     forget

i.e., all of the choices which begin with ‘f’. Pressing SPC or TAB would result in ‘fo’ appearing in the echo area, since all of the choices which begin with ‘f’ continue with ‘o’. Now, typing ‘l’ followed by ‘TAB’ results in ‘foliate’ appearing in the echo area, since that is the only choice which begins with

‘fol’.
ESC C-V (echo-area-scroll-completions-window)

Scroll the completions window, if that is visible, or the “other” window if not.

9 Printing Nodes

In general, we recommend that you use TEX to format the document and print sections of it, by running tex on the Texinfo source file. However, you may wish to print out the contents of a node as a quick reference document for later use, or if you don’t have TEX installed. Info provides you with a command for doing this.

M-x print-node

Pipe the contents of the current node through the command in the environment variable INFO_PRINT_COMMAND. If the variable does not exist, the node is simply piped to lpr (on DOS/Windows, the default is to print the node to the local printer device, ‘PRN’). The value of INFO_PRINT_COMMAND may begin with the ‘>’ character, as in ‘>/dev/printer’, in which case Info treats the rest as the name of a file or a device. Instead of piping to a command, Info opens the file, writes the node contents, and closes the file, under the assumption that text written to that file will be printed by the underlying OS.

10 Miscellaneous Commands

GNU Info contains several commands which self-document GNU Info:

M-x describe-command

Read the name of an Info command in the echo area and then display a brief description of what that command does.

M-x describe-key

Read a key sequence in the echo area, and then display the name and documentation of the Info command that the key sequence invokes.

M-x describe-variable

Read the name of a variable in the echo area and then display a brief description of what the variable affects.

M-x where-is

Read the name of an Info command in the echo area, and then display a key sequence which can be typed in order to invoke that command.

C-h (get-help-window)
?

F1 (on DOS/Windows only) h, vi-like operation

Create (or Move into) the window displaying *Help*, and place a node containing a quick reference card into it. This window displays the most concise information about GNU Info available.

h (get-info-help-node)

M-h, vi-like operation Try hard to visit the node (info)Help. The Info file ‘info.texi’ distributed with GNU Info contains this node. Of course, the file must first be processed with makeinfo, and then placed into the location of your Info directory.

Here are the commands for creating a numeric argument:

C-u (universal-argument)

Start (or multiply by 4) the current numeric argument. ‘C-u’ is a good way to give a small numeric argument to cursor movement or scrolling commands; ‘C-u C-v’ scrolls the screen 4 lines, while ‘C-u C-u C-n’ moves the cursor down 16 lines. ‘C-u’ followed by digit keys sets the numeric argument to the number thus typed: C-u 1 2 0 sets the argument to 120.

M-1 (add-digit-to-numeric-arg)

1, vi-like operation

M-2 . . . M-9

2 … 9, vi-like operation

M-0

0, vi-like operation Add the digit value of the invoking key to the current numeric argument. Once Info is reading a numeric argument, you may just type the digits of the argument, without the Meta prefix. For example, you might give ‘C-l’ a numeric argument of 32 by typing:

               C-u 3 2 C-l

or

               M-3 2 C-l
M-- (add-digit-to-numeric-arg

- To make a negative argument, type -. Typing - alone makes a negative argument with a value of -1. If you continue to type digit or Meta-digit keys after -, the result is a negative number produced by those digits.

- doesn’t work when you type in the echo area, because you need to be able to insert the ‘-’ character itself; use M-- instead, if you need to specify negative arguments in the echo area.

‘C-g’ is used to abort the reading of a multi-character key sequence, to cancel lengthy operations (such as multi-file searches) and to cancel reading input in the echo area.

C-g (abort-key)

C-u, vi-like operation Cancel current operation.

The ‘q’ command of Info simply quits running Info. Under ‘--vi-keys’ (see [–vi-keys], page 4), you can also exit with ‘:q’ or ‘ZZ’.

q (quit)
C-x C-c

:q, vi-like operation ZZ, vi-like operation

Exit GNU Info.

If the operating system tells GNU Info that the screen is 60 lines tall, and it is actually only 40 lines tall, here is a way to tell Info that the operating system is correct.

M-x set-screen-height

Read a height value in the echo area and set the height of the displayed screen to that value.

On MS-DOS/MS-Windows, this command actually tries to change the dimensions of the visible screen to the value you type in the echo area.

Finally, Info provides a convenient way to display footnotes which might be associated with the current node that you are viewing:

ESC C-f (show-footnotes)

Show the footnotes (if any) associated with the current node in another window. You can have Info automatically display the footnotes associated with a node when the node is selected by setting the variable automatic-footnotes. See

          Chapter 11 [automatic-footnotes], page 26.

11 Manipulating Variables

GNU Info contains several variables whose values are looked at by various Info commands. You can change the values of these variables, and thus change the behavior of Info to more closely match your environment and Info file reading manner.

There are two ways to set the value of a variable: interactively, using the set-variable command described below, or in the #var section of the .infokey file. See Chapter 12 [Custom Key Bindings], page 29.

M-x set-variable

Read the name of a variable, and the value for it, in the echo area and then set the variable to that value. Completion is available when reading the variable name (see Section 8.3 [The Echo Area], page 19); often, completion is available when reading the value to give to the variable, but that depends on the variable itself. If a variable does not supply multiple choices to complete over, it expects a numeric value.

M-x describe-variable

Read the name of a variable in the echo area and then display a brief description of what the variable affects.

Here is a list of the variables that you can set in Info.

automatic-footnotes

When set to On, footnotes appear and disappear automatically; else, they appear at the bottom of the node text. This variable is Off by default. When a node is selected, a window containing the footnotes which appear in that node is created, and the footnotes are displayed within the new window. The window that Info creates to contain the footnotes is called ‘*Footnotes*’. If a node is selected which contains no footnotes, and a ‘*Footnotes*’ window is on the screen, the ‘*Footnotes*’ window is deleted. Footnote windows created in this fashion are not automatically tiled so that they can use as little of the display as is possible.

automatic-tiling

When set to On, creating or deleting a window resizes other windows. This variable is Off by default. Normally, typing ‘C-x 2’ divides the current window into two equal parts. When automatic-tiling is set to On, all of the windows are resized automatically, keeping an equal number of lines visible in each window. There are exceptions to the automatic tiling; specifically, the windows ‘*Completions*’ and ‘*Footnotes*’ are not resized through automatic tiling; they remain their original size.

cursor-movement-scrolls

Normally, cursor movement commands (see Chapter 3 [Cursor Commands], page 6) stop when top or bottom of a node is reached. When this variable is set to On, cursor movement commands act as scrolling ones and their behavior is controlled by the scroll-behavior variable (see below).

errors-ring-bell

When set to On, errors cause the bell to ring. The default setting of this variable is On.

gc-compressed-files

When set to On, Info garbage collects files which had to be uncompressed. The default value of this variable is Off. Whenever a node is visited in Info, the Info file containing that node is read into core, and Info reads information about the tags and nodes contained in that file. Once the tags information is read by Info, it is never forgotten. However, the actual text of the nodes does not need to remain in core unless a particular Info window needs it. For non-compressed files, the text of the nodes does not remain in core when it is no longer in use. But de-compressing a file can be a time consuming operation, and so Info tries hard not to do it twice. gc-compressed-files tells Info it is okay to garbage collect the text of the nodes of a file which was compressed on disk.

ISO-Latin

When set to On, Info accepts and displays ISO Latin characters. By default, Info assumes an ASCII character set. ISO-Latin tells Info that it is running in an environment where the European standard character set is in use, and allows you to input such characters to Info, as well as display them.

scroll-behavior
scroll-behaviour

Control what happens when forward scrolling is requested at the end of a node, or when backward scrolling is requested at the beginning of a node. The default value for this variable is Continuous. There are three possible values for this variable:

Continuous

Try to get the first item in this node’s menu, or failing that, the ‘Next’ node, or failing that, the ‘Next’ of the ‘Up’. This behavior is identical to using the ‘]’ (global-next-node) and ‘[’ (globalprev-node) commands.

scroll-last-node

Control what happens when a scrolling command is issued at the end of the last node. Possible values are:

Stop Do not scroll. Display the ‘No more nodes within this document.’ message. This is the default.

Scroll Scroll as usual. Since the last node is usually an index, this means that the very first node from the menu will be selected.

Top Go to the top node of this document.

This variable is in effect only if scroll-behaviour is set to

          Continuous.

Notice that the default behavior has changed in version 4.12. Previous versions behaved as if scroll-last-node=Scroll was set.

This behavior was counter-intuitive, therefore since version 4.12 the default is to stop at the last node.

Next Only Only try to get the ‘Next’ node.

Page Only Simply give up, changing nothing. If scroll-behavior is Page Only, no scrolling command can change the node that is being viewed.

This variable normally affects only scrolling commands. See [cursor-movement-scrolls], page 26, for information on how to widen its scope.

The two names, scroll-behavior and scroll-behaviour, are a historical

accident. They are merely synonyms.

scroll-step

The number of lines to scroll when the cursor moves out of the window. Scrolling happens automatically if the cursor has moved out of the visible portion of the node text when it is time to display. Usually the scrolling is done so as to put the cursor on the center line of the current window. However, if the variable scroll-step has a nonzero value, Info attempts to scroll the node text by that many lines; if that is enough to bring the cursor back into the window, that is what is done. The default value of this variable is 0, thus placing the cursor (and the text it is attached to) in the center of the window. Setting this variable to 1 causes a kind of “smooth scrolling” which some people prefer.

show-index-match

When set to On, the portion of the matched search string is highlighted in the message which explains where the matched search string was found. The default value of this variable is On. When Info displays the location where an index match was found, (see Chapter 6 [next-index-match], page 14), the portion of the string that you had typed is highlighted by displaying it in the inverse case from its surrounding characters.

visible-bell

When set to On, GNU Info attempts to flash the screen instead of ringing the bell. This variable is Off by default. Of course, Info can only flash the screen if the terminal allows it; in the case that the terminal does not allow it, the setting of this variable has no effect. However, you can make Info perform quietly by setting the errors-ring-bell variable to Off.

12 Customizing Key Bindings and Variables

GNU Info provides a way to define arbitrary key-to-command bindings and variable settings, overriding the defaults described in this document. (The ‘--vi-keys’ option rebinds many keys at once; see [–vi-keys], page 4.)

On startup, GNU Info looks for a configuration file in the invoker’s HOME directory called ‘.info’^{1}. If it is present, and appears to contain Info configuration data, and was created with the current version of the infokey command, then Info adopts the key bindings and variable settings contained therein.

The ‘.info’ file contains compact, non-textual data for reasons of efficiency and because its design was lifted wholesale from the GNU Less program, which also does it that way. It must be created by compiling a textual source file using the infokey command.

12.1 Invoking infokey

infokey compiles a source file (‘$HOME/.infokey’^{2} by default) containing Info customizations into a binary format (‘$HOME/.info’ by default). GNU Info reads the binary file at startup to override the default key bindings and variable definitions. Synopsis:

infokey [option...] [input-file]

Besides the standard ‘--help’ and ‘--version’, the only option is ‘--output file’. This tells infokey to write the binary data to file instead of ‘$HOME/.info’.

12.2 infokey source format

The format of the source file read by infokey is most easily illustrated by example. For instance, here is a sample ‘.infokey’ source file suitable for aficionados of vi or less:

#info
j       next-line
k       prev-line
l       forward-char
h       backward-char
\kd     next-line
\ku     prev-line
\kr     forward-char
\kl     backward-char
\       scroll-forward
\kD     scroll-forward-page-only
b       scroll-backward
\kU     scroll-backward-page-only
g       beginning-of-node
\kh     beginning-of-node
G       end-of-node

1 Due to the limitations of DOS filesystems, the MS-DOS version of Info looks for a file ‘_info’ instead.

If the HOME variable is not defined, Info additionally looks in the current directory.

2 This file is named ‘_infokey’ in the MS-DOS version, and is looked for in the current directory if HOME

is undefined.

\ke     end-of-node
\t      select-reference-this-line
-       history-node
n       next-node
p       prev-node
u       up-node
t       top-node
d       dir-node
#var
scroll-step=1

The source file consists of one or more sections. Each section starts with a line that identifies the type of section. Possible sections are:

#info Key bindings for Info windows. The start of this section is indicated by a line containing just #info by itself. If this is the first section in the source file, the #info line can be omitted. The rest of this section consists of lines of the form:

     string whitespace action [ whitespace [ # comment ] ] newline

Whitespace is any sequence of one or more spaces and/or tabs. Comment is any sequence of any characters, excluding newline. string is the key sequence which invokes the action. action is the name of an Info command. The characters in string are interpreted literally or prefixed by a caret (^) to indicate a control character. A backslash followed by certain characters specifies input keystrokes as follows:

\b Backspace

\e Escape (ESC)

\n Newline

\r Return

\t Tab

\ku Up arrow

\kd Down arrow

\kl Left arrow

\kr Right arrow

\kU Page Up

\kD Page Down

\kh HOME

\ke  END

\kx Delete (DEL)

\mx Meta-x where x is any character as described above.

Backslash followed by any other character indicates that character is to be taken literally. Characters which must be preceded by a backslash include caret, space, tab, and backslash itself.

#echo-area

Key bindings for the echo area. The start of this section is indicated by a line containing just #echo-area by itself. The rest of this section has a syntax identical to that for the key definitions for the Info area, described above.

#var Variable initializations. The start of this section is indicated by a line containing just #var by itself. Following this line is a list of variable assignments, one per line. Each line consists of a variable name (See Chapter 11 [Variables], page 26,) followed by = followed by a value. There may be no white space between the variable name and the =, and all characters following the =, including white space, are included in the value.

Blank lines and lines starting with # are ignored, except for the special section header lines.

Key bindings defined in the ‘.info’ file take precedence over GNU Info’s default key bindings, whether or not ‘--vi-keys’ is used. A default key binding may be disabled by overriding it in the ‘.info’ file with the action invalid. In addition, all default key bindings can be disabled by adding this line anywhere in the relevant section:

     #stop

This will cause GNU Info to ignore all the default key commands for that section.

Beware: #stop can be dangerous. Since it disables all default key bindings, you must supply enough new key bindings to enable all necessary actions. Failure to bind any key to the quit command, for example, can lead to frustration.

The order in which key bindings are defined in the ‘.info’ file is not important, except that the command summary produced by the get-help-window command only displays the first key that is bound to each command.

Appendix A: Index 32

Appendix A Index

’

', vi-like operation . . . . . . . . . . . . . . . . . . . . . . . . . 11

,

, . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

-

- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 --subnodes, command line option. . . . . . . . . . . . . . . 4

.

.info . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 .infokey . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 .infokey source format . . . . . . . . . . . . . . . . . . . . . . . . . 29

/

/ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

<

< . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

>

> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

?

?, in Info windows. . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 ?, in the echo area . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 ?, vi-like operation . . . . . . . . . . . . . . . . . . . . . . . . . 14

[

[ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

]

] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

info file (MS-DOS) . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 infokey file (MS-DOS) . . . . . . . . . . . . . . . . . . . . . . . . 29

0

0 ... 9, vi-like operation . . . . . . . . . . . . . . . . . . . 24

0, in Info windows. . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

1

1 ... 9, in Info windows . . . . . . . . . . . . . . . . . . . . . 16

A

abort-key . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

add-digit-to-numeric-arg . . . . . . . . . . . . . . . . . . . 24

ANSI escape sequences in documents. . . . . . . . . . . . 3 Apropos, in Info files. . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 arguments, command line . . . . . . . . . . . . . . . . . . . . . . . 2 arguments, negative . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 automatic-footnotes . . . . . . . . . . . . . . . . . . . . . . . . . 26 automatic-tiling. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26

B

b, in Info windows. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 b, vi-like operation . . . . . . . . . . . . . . . . . . . . . . . . . . 9 BackTab, in Info windows. . . . . . . . . . . . . . . . . . . . . 17 BackTab, in the echo area . . . . . . . . . . . . . . . . . . . . 21 backward-char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 backward-word . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 beginning-of-line. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 beginning-of-node. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 BS (backspace) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

C

C-a, in Info windows . . . . . . . . . . . . . . . . . . . . . . . . . . 6 C-a, in the echo area. . . . . . . . . . . . . . . . . . . . . . . . . 20 C-b, in Info windows . . . . . . . . . . . . . . . . . . . . . . . . . . 6 C-b, in the echo area. . . . . . . . . . . . . . . . . . . . . . . . . 20 C-b, vi-like operation. . . . . . . . . . . . . . . . . . . . . . . . 9 C-CENTER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 C-d, in the echo area. . . . . . . . . . . . . . . . . . . . . . . . . 20 C-d, vi-like operation. . . . . . . . . . . . . . . . . . . . . . . . 9 C-e, in Info windows . . . . . . . . . . . . . . . . . . . . . . . . . . 6 C-e, in the echo area. . . . . . . . . . . . . . . . . . . . . . . . . 20 C-e, vi-like operation. . . . . . . . . . . . . . . . . . . . . . . . 9 C-End . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 C-f, in Info windows . . . . . . . . . . . . . . . . . . . . . . . . . . 6 C-f, in the echo area. . . . . . . . . . . . . . . . . . . . . . . . . 20 C-f, vi-like operation. . . . . . . . . . . . . . . . . . . . . . . . 8 C-g, in Info windows . . . . . . . . . . . . . . . . . . . . . . . . . 25 C-g, in the echo area. . . . . . . . . . . . . . . . . . . . . . . . . 20 C-h. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 C-Home . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 C-k, in the echo area. . . . . . . . . . . . . . . . . . . . . . . . . 21 C-k, vi-like operation. . . . . . . . . . . . . . . . . . . . . . . . 9 C-l. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 C-LEFT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 C-LEFT, in the echo area . . . . . . . . . . . . . . . . . . . . . 20 C-n . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 C-n, vi-like operation. . . . . . . . . . . . . . . . . . . . . . . . 9 C-NEXT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 C-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 C-p, vi-like operation. . . . . . . . . . . . . . . . . . . . . . . . 9 C-PgDn . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 C-PgUp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 C-PREVIOUS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 C-q, in the echo area. . . . . . . . . . . . . . . . . . . . . . . . . 20 C-r. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 C-RIGHT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 C-RIGHT, in the echo area . . . . . . . . . . . . . . . . . . . . 20 C-s. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 C-t, in the echo area. . . . . . . . . . . . . . . . . . . . . . . . . 21 C-u. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

C-u cancels typeahead, vi-like operation. . . 25
C-u, in the echo area, vi-like operation . . . 20

C-u, vi-like operation. . . . . . . . . . . . . . . . . . . . . . . . 9 C-UP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 C-v . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

C-v, in the echo area, vi-like operation . . . 20

C-w. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 C-x ^. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 C-x 0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 C-x 1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 C-x 2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 C-x b . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 C-x C-b . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 C-x C-c . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 C-x C-f . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 C-x DEL, in the echo area . . . . . . . . . . . . . . . . . . . . 21 C-x g, vi-like operation. . . . . . . . . . . . . . . . . . . . . 12 C-x k . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 C-x n . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 C-x N . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 C-x n, vi-like operation. . . . . . . . . . . . . . . . . . . . . 11 C-x o . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 C-x r, vi-like operation. . . . . . . . . . . . . . . . . . . . . 17 C-x t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 C-x u, vi-like operation. . . . . . . . . . . . . . . . . . . . . 11 C-y, in the echo area. . . . . . . . . . . . . . . . . . . . . . . . . 21 C-y, vi-like operation. . . . . . . . . . . . . . . . . . . . . . . . 9 cancelling the current operation. . . . . . . . . . . . . . . . 25 cancelling typeahead. . . . . . . . . . . . . . . . . . . . . . . . . . . 25 case-sensitive search . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 case-sensitivity, and search. . . . . . . . . . . . . . . . . . . . . 15 colors in documents. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 command line options. . . . . . . . . . . . . . . . . . . . . . . . . . . 2 command-line options, how to find . . . . . . . . . . . . . . 4 commands, describing . . . . . . . . . . . . . . . . . . . . . . . . . 24 completion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 compressed Info files . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 cursor, moving. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 customizing key bindings . . . . . . . . . . . . . . . . . . . . . . 29

D

d . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 d, vi-like operation . . . . . . . . . . . . . . . . . . . . . . . . . . 9 default key bindings, overriding . . . . . . . . . . . . . . . . 29 DEL, in Info windows . . . . . . . . . . . . . . . . . . . . . . . . . . 8 DEL, in the echo area. . . . . . . . . . . . . . . . . . . . . . . . . 20 delete-window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 describe-command. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 describe-key . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 describe-variable. . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 dir-node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 directory path . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 DOWN (an arrow key). . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 DOWN, vi-like operation. . . . . . . . . . . . . . . . . . . . . . . 9 down-line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

E

e, in Info windows. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 echo area . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 echo-area-abort . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 echo-area-backward . . . . . . . . . . . . . . . . . . . . . . . . . . 20

                                                echo-area-backward-kill-line . . . . . . . . . . . . . . . 21
                                                echo-area-backward-kill-word . . . . . . . . . . . . . . . 21
                                                echo-area-backward-word. . . . . . . . . . . . . . . . . . . . . 20

echo-area-beg-of-line . . . . . . . . . . . . . . . . . . . . . . . 20 echo-area-complete . . . . . . . . . . . . . . . . . . . . . . . . . . 21 echo-area-delete. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 echo-area-end-of-line . . . . . . . . . . . . . . . . . . . . . . . 20 echo-area-forward. . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 echo-area-forward-word . . . . . . . . . . . . . . . . . . . . . . 20 echo-area-insert. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 echo-area-kill-line . . . . . . . . . . . . . . . . . . . . . . . . . 21 echo-area-kill-word . . . . . . . . . . . . . . . . . . . . . . . . . 21 echo-area-newline. . . . . . . . . . . . . . . . . . . . . . . . . . . . 20

                                                echo-area-possible-completions. . . . . . . . . . . . . 22
                                                echo-area-quoted-insert. . . . . . . . . . . . . . . . . . . . . 20

echo-area-rubout. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20

                                                echo-area-scroll-completions-window . . . . . . . 22

echo-area-tab-insert . . . . . . . . . . . . . . . . . . . . . . . . 21

                                                echo-area-transpose-chars . . . . . . . . . . . . . . . . . . 21

echo-area-yank . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 echo-area-yank-pop . . . . . . . . . . . . . . . . . . . . . . . . . . 21 End . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 end-of-line. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 end-of-node. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 errors-ring-bell. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 ESC C-f . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 ESC C-v, in Info windows . . . . . . . . . . . . . . . . . . . . . 19 ESC C-v, in the echo area . . . . . . . . . . . . . . . . . . . . 22

F

f . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 f, vi-like operation . . . . . . . . . . . . . . . . . . . . . . . . . . 8 F1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 file names, relative . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 file, outputting to. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 files, compressed . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 find-menu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 finding the Invocation node . . . . . . . . . . . . . . . . . . . . 12 first-node. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 footnotes, displaying . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 format of .infokey source. . . . . . . . . . . . . . . . . . . . . . . 29 forward-char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 forward-word . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 functions, describing . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

G

g . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 G . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 g, vi-like operation . . . . . . . . . . . . . . . . . . . . . . . . . 12 G, vi-like operation . . . . . . . . . . . . . . . . . . . . . . . . . 12 gc-compressed-files . . . . . . . . . . . . . . . . . . . . . . . . . 27 get-help-window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 get-info-help-node . . . . . . . . . . . . . . . . . . . . . . . . . . 24 global-next-node. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 global-prev-node. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 goto-invocation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 goto-node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 grow-window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

H

h . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 h, vi-like operation . . . . . . . . . . . . . . . . . . . . . . . . . 24 history-node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 Home. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

I

i . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 I . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 incremental search . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 index search, selecting from the command line… 3 index, searching . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 index-apropos . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 index-search . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Info file, selecting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Info files, compressed . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Info files, relative . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Info files, searching all indices . . . . . . . . . . . . . . . . . . . 2 Info manual location . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Info, invoking. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 INFO PRINT COMMAND, environment variable

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 infokey . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 infokey source format . . . . . . . . . . . . . . . . . . . . . . . . . . 29 infokey, invoking. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 invocation description, how to find . . . . . . . . . . . . . . 4 invoking Info . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 invoking infokey . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 isearch-backward. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 isearch-forward . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 ISO Latin characters . . . . . . . . . . . . . . . . . . . . . . . . . . 27 ISO-Latin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27

K

k, vi-like operation . . . . . . . . . . . . . . . . . . . . . . . . . . 9 keep-one-window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 key bindings, customizing . . . . . . . . . . . . . . . . . . . . . . 29 keys, describing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 keystrokes, recording. . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 kill-node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

L

l . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 last-menu-item . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 last-node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 LEFT (an arrow key). . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 LEFT, in the echo area . . . . . . . . . . . . . . . . . . . . . . . 20 Less-like key bindings . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LFD, vi-like operation. . . . . . . . . . . . . . . . . . . . . . . . 9 list-visited-nodes . . . . . . . . . . . . . . . . . . . . . . . . . . 13 local printer device . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

M

m . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 M-$, vi-like operation. . . . . . . . . . . . . . . . . . . . . . . 20 M--. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 M-< . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 M-> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 M-0 ... M-9 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

M-0, in the echo area, vi-like operation . . . 20

M-0, vi-like operation. . . . . . . . . . . . . . . . . . . . . . . 17

M-1 ... M-9, vi-like operation . . . . . . . . . . . . . . 16

M-b, in Info windows . . . . . . . . . . . . . . . . . . . . . . . . . . 7 M-b, in the echo area. . . . . . . . . . . . . . . . . . . . . . . . . 20 M-b, vi-like operation. . . . . . . . . . . . . . . . . . . . . . . . 7 M-BS, in the echo area . . . . . . . . . . . . . . . . . . . . . . . 21 M-d, in the echo area. . . . . . . . . . . . . . . . . . . . . . . . . 21 M-d, vi-like operation. . . . . . . . . . . . . . . . . . . . . . . 11 M-DEL, in the echo area . . . . . . . . . . . . . . . . . . . . . . 21 M-f, in Info windows . . . . . . . . . . . . . . . . . . . . . . . . . . 6 M-f, in the echo area. . . . . . . . . . . . . . . . . . . . . . . . . 20 M-f, vi-like operation. . . . . . . . . . . . . . . . . . . . . . . 17 M-g, vi-like operation. . . . . . . . . . . . . . . . . . . . . . . 17

M-h, in the echo area, vi-like operation . . . 20

M-h, vi-like operation. . . . . . . . . . . . . . . . . . . . . . . 24

M-l, in the echo area, vi-like operation . . . 20

M-r . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

M-SPC, vi-like operation . . . . . . . . . . . . . . . . . . . . . 8

M-t, vi-like operation. . . . . . . . . . . . . . . . . . . . . . . 11 M-TAB, in Info windows . . . . . . . . . . . . . . . . . . . . . . . 17 M-TAB, in the echo area . . . . . . . . . . . . . . . . . . . . . . 21 M-v . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

M-w, in the echo area, vi-like operation . . . 20
M-x, in the echo area, vi-like operation . . . 20
M-X, in the echo area, vi-like operation . . . 21

M-y, in the echo area. . . . . . . . . . . . . . . . . . . . . . . . . 21 menu, following . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 menu, following, from inside Info. . . . . . . . . . . . . . . 13 menu-digit. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 menu-item . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 menu-sequence . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 move-to-next-xref. . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 move-to-prev-xref. . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 move-to-window-line . . . . . . . . . . . . . . . . . . . . . . . . . . 7 moving the cursor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

N

n . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 n, vi-like operation . . . . . . . . . . . . . . . . . . . . . . . . . 14 negative arguments . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 NEXT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 next-index-match. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 next-line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 next-node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 next-window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 node, selecting from the command line . . . . . . . . . . 3 nodes, selection of . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 numeric arguments . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 numeric arguments, negative . . . . . . . . . . . . . . . . . . . 25

O

O . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 online help, using Info as. . . . . . . . . . . . . . . . . . . . . . . . 3 options, command line . . . . . . . . . . . . . . . . . . . . . . . . . . 2 outputting to a file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 overriding default key bindings. . . . . . . . . . . . . . . . . 29

P

p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 PageDown . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 PageUp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 prev-line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 prev-node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 prev-window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 PREVIOUS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 print-node. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 printing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

printing characters, in the echo area. . . . . . . 21

printing nodes to the local printer. . . . . . . . . . . . . . 23

Q

q . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 quit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 quitting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

R

r . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 R . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 redraw-display . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 regular expression search . . . . . . . . . . . . . . . . . . . . . . 14 relative Info file names . . . . . . . . . . . . . . . . . . . . . . . . . . 2 remembering user keystrokes . . . . . . . . . . . . . . . . . . . . 2 repeated search . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 replaying recorded keystrokes . . . . . . . . . . . . . . . . . . . 4 RET, in Info windows . . . . . . . . . . . . . . . . . . . . . . . . . 17 RET, in the echo area. . . . . . . . . . . . . . . . . . . . . . . . . 20 RET, vi-like operation. . . . . . . . . . . . . . . . . . . . . . . . 9 RIGHT (an arrow key) . . . . . . . . . . . . . . . . . . . . . . . . . . 6 RIGHT, in the echo area . . . . . . . . . . . . . . . . . . . . . . 20

S

s . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 S . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 screen, changing the height of . . . . . . . . . . . . . . . . . . 25 scroll-backward . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

scroll-backward-page-only . . . . . . . . . . . . . . . . . . . 9
scroll-backward-page-only-set-window . . . . . . . 9

scroll-behavior . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 scroll-behaviour. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 scroll-forward . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

scroll-forward-page-only. . . . . . . . . . . . . . . . . . . . . 8
scroll-forward-page-only-set-window . . . . . . . . 8
scroll-half-screen-down . . . . . . . . . . . . . . . . . . . . . . 9

scroll-half-screen-up . . . . . . . . . . . . . . . . . . . . . . . . 9 scroll-last-node. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 scroll-other-window . . . . . . . . . . . . . . . . . . . . . . . . . 19 scroll-step . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 scrolling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 scrolling through node structure . . . . . . . . . . . . . . . . 9 search . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 search, and case-sensitivity . . . . . . . . . . . . . . . . . . . . 15 search, case-sensitive. . . . . . . . . . . . . . . . . . . . . . . . . . . 14 search-backward . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

search-case-sensitively. . . . . . . . . . . . . . . . . . . . . 14

search-next . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 search-previous . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 searching . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 Searching all indices . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 searching, in the indices . . . . . . . . . . . . . . . . . . . . . . . 15

select-reference-this-line . . . . . . . . . . . . . . . . . 17

select-visited-node . . . . . . . . . . . . . . . . . . . . . . . . . 13 set-screen-height. . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 set-variable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26

Shift-TAB, in Info windows . . . . . . . . . . . . . . . . . . 17
Shift-TAB, in the echo area. . . . . . . . . . . . . . . . . . 21

show-footnotes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 show-index-match. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 SPC, in Info windows . . . . . . . . . . . . . . . . . . . . . . . . . . 8 SPC, in the echo area. . . . . . . . . . . . . . . . . . . . . . . . . 21 speech synthesizers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 split-window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

T

t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 TAB, in Info windows . . . . . . . . . . . . . . . . . . . . . . . . . 17 TAB, in the echo area. . . . . . . . . . . . . . . . . . . . . . . . . 21 tile-windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 tiling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 toggle-regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 toggle-wrap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 top-node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

U

u . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 u, vi-like operation . . . . . . . . . . . . . . . . . . . . . . . . . . 9 universal-argument . . . . . . . . . . . . . . . . . . . . . . . . . . 24 UP (an arrow key) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 UP, vi-like operation . . . . . . . . . . . . . . . . . . . . . . . . . 9 up-line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 up-node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

V

variables, describing . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 variables, setting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 version information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 vi-like key bindings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 view-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 visible-bell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

W

w, vi-like operation . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Where is an Info manual? . . . . . . . . . . . . . . . . . . . . . . . 4 where-is . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 windows, creating . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 windows, deleting. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 windows, manipulating . . . . . . . . . . . . . . . . . . . . . . . . 18 windows, selecting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

X

xref-item . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

Y

y, vi-like operation . . . . . . . . . . . . . . . . . . . . . . . . . . 9

Z

z, vi-like operation . . . . . . . . . . . . . . . . . . . . . . . . . . 8 ZZ, vi-like operation . . . . . . . . . . . . . . . . . . . . . . . . 25